OpenStudy (jenniferjuice):

Evaluate the following expression using the values given. Find 4x - y - 2z if x = -2, y = 3, and z = -2

OpenStudy (jenniferjuice):

-15 15 -7 7

OpenStudy (anonymous):

We just need to plug in the given values for the variables. So, plugging in x=-2, y=3 and z=-2 we get: \[4(-2) -3 -2(-2)\]Simplifying this we get: \[-8 -3 +4 = -11 +4=-7\]